OpenStudy (anonymous):

i need a step by step tutorial on how to do 3-variable equations... i learned how to at school and i got a take home test and every answer i do comes out wrong. anyone want to help?

OpenStudy (anonymous):

you mean 3 variable 3 equations system?

OpenStudy (anonymous):

im not sure what theyre called. it looks like this: x+y+z=1 2x+y=4 x+z=6 ( i made that up)

OpenStudy (anonymous):

x+y+z=1 2x+y=4 x+z=6 reduce that to two equation z=-x+6 plug in first equation for z x+y+(-x+6)=1 2x+y=4 now you got two equation and two unknown which is easier to solve

OpenStudy (anonymous):

so if my question was x-2y=1 x+3y+1z=0 2x-2x=18 would my 2 new equations be 5y+1z=-1 and 8x-y=11 ?

OpenStudy (anonymous):

show me what you did, like what did you subsitute

OpenStudy (anonymous):

i did x-2y+0z=1 x+3y+1z=0 but then i had to multiply by -1 in order to substitute the x so then i got -x+2y-0z=-1 x+3y+1z=0 which is when i got my first new equation 5y+1z=-1 for the second one i did x+3y+1z=0 2x+0y-2z=18 but i had to multiply by -2 to get rid of the x so i got -2x-6x-2z=0 2x+0y-2z=18 which is when i got my 2nd new equation -6y-4z=18